### v1

# The Modularity of an Abelian Variety

Abstract.
We introduce the concept of the modularity of an abelian variety defined over the rational number field extending the modularity of an elliptic curve. We discuss the modularity of an abelian variety over $\mathbb{Q}$. We conjecture that a simple abelian variety over $\mathbb{Q}$ is modular. The detailed article will appear later.
Subject Classification: Primary 14Kxx, 14G35, 11F46, 11F80
Keywords and phrases: modularity, abelian varieties, Siegel modular variety, Siegel modular forms, Galois
representation, zeta functions
1. The Modularity of an Elliptic Curve
We set $\Gamma_{1}:=SL(2,\mathbb{Z})$. For a positive integer $N$, we let $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$ be the congruence subgroups of $\Gamma_{1}$ such that $\Gamma(N)\subset\Gamma_{1}(N)\subset\Gamma_{0}(N)\subset\Gamma_{1}.$ We refer to [6, pp. 13–14, p. 21] for the precise definitions and properties of $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$. Let $\mathbb{H}_{1}$ be the Poincaré upper half plance. The quotient
$Y_{1}(N):=\Gamma_{1}(N)\backslash\mathbb{H}_{1}\ ({\rm resp}.\ Y_{0}(N):=\Gamma_{0}(N)\backslash\mathbb{H}_{1})$
be the complex manifold which has a natural model $Y_{1}(N)/\mathbb{Q}$ (resp. $Y_{0}(N)/\mathbb{Q}$). We let $X_{1}(N)$ (resp. $X_{0}(N)$) be the smooth projective curve which contains $Y_{1}(N)$ (resp. $Y_{0}(N)$) as a dense Zariski open subset (cf. see [6, pp. 45–60]).
Let $S_{k}(N)$ be the space of cusp forms of weight $k\geq 1$ and level $N\geq 1$. Here $k$ and $N$ be positive integers. We recall that if $f\in S_{k}(N)$, it satisfies the following properties : (C1) $f((a\tau+b)(c\tau+d)^{-1})=(c\tau+d)^{k}f(\tau)$ for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}(N)$ and $\tau\in\mathbb{H}_{1}$; (C2) $|f(\tau)|^{2}({\rm Im}\,\tau)^{k}$ is bounded in $\mathbb{H}_{1}$ and (C3) the Fourier expansion of $f(\tau)$ is given by
$f(\tau)=\sum_{n=1}^{\infty}a_{n}(f)\,q^{n},\quad{\rm where}\ q=e^{2\pi i\tau}.$
We define the $L$-series of $f\in S_{k}(N)$ to be
$L(f,s):=\sum_{n=1}^{\infty}a_{n}(f)\,n^{-s}.$
For each prime $p\!\not|N$, we recall that the Hecke operator $T_{p}:S_{k}(N)\longrightarrow S_{k}(N)$ is defined by
$(T_{p}f)(\tau)=p^{-1}\sum_{i=0}^{p-1}f\left(\frac{\tau+i}{p}\right)+p^{k-1}f\left(\frac{ap\tau+b}{cp\tau+d}\right),\quad f\in S_{k}(N)$
for any $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}$ with $c\equiv 0\,({\rm mod}\,N)$ and $d\equiv p\,({\rm mod}\,N)$. We refer to [2, p. 844] or [6, pp. 170–171] for more detail. The Hecke operators $T_{p}\,(p\!\!\not|N)$ can be simultaneously diagonalized on $S_{k}(N)$ and a simultaneous eigenvector a Hecke eigenform or simply an eigenform.
Let $\lambda$ be a place of the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ lying over a rational integer $\ell$ and ¯${\bar{\mathbb{Q}}}_{\lambda}$ denote the algebraic closure of $\mathbb{Q}_{\ell}$ via $\lambda$. Let $G_{\mathbb{Q}}:={\rm Gal}({\bar{\mathbb{Q}}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$. It is well known that if $f\in S_{k}(N)$ is a normalized eigenform with $a_{1}(f)=1$, then there exists a unique continuous irreducible Galois representation
$\rho_{f,\lambda}:G_{\mathbb{Q}}\longrightarrow GL(2,{\bar{\mathbb{Q}}}_{\lambda})$
such that $\rho_{f,\lambda}$ is unramified at $p$ for all primes $p\!\!\not|\,\ell N$ and
${\rm Tr}\left(\rho_{f,\lambda}({\rm Frob}_{p})\right)=a_{p}(f)\quad{\rm for\ any\ prime}\ p\!\!\not|\,\ell N.$
The existence of $\rho_{f,\lambda}$ is due to Shimura if $k=2$ [17], due to Deligne if $k>2$ [4] and due to Deligne and Serre if $k=1$ [5]. We see that $\rho_{f,\lambda}$ is odd in the sense that $\det\rho_{f,\lambda}$ of complex conjugation is -1. Moreover $\rho_{f,\lambda}$ is potentially semi-stable at $\ell$ in the sense of Fontaine [9].
We may choose a conjugate of $\rho_{f,\lambda}$ which is valued in $GL(2,{\mathcal{O}}_{\bar{\mathbb{Q}}_{\lambda}})$ and reducing modulo the maximal ideal and semi-simplyfing yields an irreducible continuous representation
${\overline{\rho}}_{f,\lambda}:G_{\mathbb{Q}}\longrightarrow GL(2,\overline{\mathbb{F}}_{\ell})$
which, up to isomorphism, does not depend on the choice of conjugate of $\rho_{f,\lambda}$.
Definition 1.1.
Let
$\rho:G_{\mathbb{Q}}\longrightarrow GL(2,\bar{\mathbb{Q}}_{\ell})$
be an irreducible continuous Galois representation which is unramified outside finitely many primes and for which the restriction of $\rho$ to a decomposition group at $\ell$ is potentially semi-stable at $\ell$ in the sense of Fontaine. Then $\rho$ is called ${\sf modular}$ if $\rho$ is equivalent to $\rho_{f,\lambda}$ (denoted $\rho\sim\rho_{f,\lambda}$) for some normalized eigenform $f$ and some place $\lambda|\ell.$
Definition 1.2.
Let
${\overline{\rho}}:G_{\mathbb{Q}}\longrightarrow GL(2,\overline{\mathbb{F}}_{\ell})$
be a two-dimensional irreducible continuous representation of $G_{\mathbb{Q}}$. Then ${\overline{\rho}}$ is called ${\sf modular}$ if ${\overline{\rho}}\sim{\overline{\rho}}_{f,\lambda}$ for some normalized eigenform $f$ and some place $\lambda|\ell.$
Let $E$ be an elliptic curve over $\mathbb{Q}$. We define
$a_{p}(E):=p+1-|E({\mathbb{F}}_{p})|\quad{\rm for\ a\ prime}\ p.$
The $L$-function $L(E,s)$ of $E$ is defined by the product of the local $L$-factors
$L(E,s):=\prod_{p|D}\left(\frac{1}{1-a_{p}(E)p^{-s}}\right)\prod_{p\not|D}\left(\frac{1}{1-a_{p}(E)\,p^{-s}+p^{1-2s}}\right).$
Then $L(E,s)$ converges absolutely for ${\rm Re}\,s>\frac{3}{2}$ and extends to an entire function by [2].
Definition 1.3.
An elliptic curve $E$ over $\mathbb{Q}$ is called ${\sf modular}$ if there exists a Hecke eigenform $f\in S_{2}(N)$ such that
$L(E,s)=L(f,s).$
Let $E$ be an elliptic curve over $\mathbb{Q}$ with its conductor $N(E)$. Let
$\rho_{E,\ell}:G_{\mathbb{Q}}\longrightarrow GL(2,\bar{\mathbb{Q}}_{\ell})$
be the $\ell$-adic representation of $G_{\mathbb{Q}}$ with the Tate module $T_{\ell}(E)$ as its representation space. Let
$J_{1}(N):=\Omega^{1}(X_{1}(N))^{\vee}/H_{1}(X_{1}(N),\mathbb{Z})\cong S_{2}(\Gamma_{1}(N))^{\vee}/H_{1}(X_{1}(N),\mathbb{Z})$
be the Jacobian variety of the modular curve $X_{1}(N)$. Here $\Omega^{1}(X_{1}(N))$ denotes the complex vector space of holomorphic 1-forms on $X_{1}(N)$ and $W^{\vee}$ denotes the dual space of a complex vector space $W$. It is known that the following statements are equivalent : (a) $E$ is modular. (b) There is a non-constant holomorphic mapping $X_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
integer $N$. (c) There is a non-constant holomorphic mapping $J_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
integer $N$. (d) $\rho_{E,\ell}$ is modular for a prime $\ell$.
The above statements have been called the Taniyama-Shimura conjecture. We refer to [18] for the historical story of this conjecture. The implication (a) $\Longrightarrow$ (b) follows from a construction of Shimura [17] and a theorem of Faltings [7]. The implication (b) $\Longrightarrow$ (d) is due to Mazur [14]. The implication (d) $\Longrightarrow$ (a) follows from a theorem of Carayol [3]. The implication (c) $\Longrightarrow$ (b) is obvious. Wiles [21, 22] proved that a semistable eliptic curve over $\mathbb{Q}$ is modular by proving the statement (d). Thereafter Breuil, Conrad, Diamond and Taylor [2] proved that every eliptic curve over $\mathbb{Q}$ is modular.
Serre [16] conjectured the following : Serre’s Modularity Conjecture : Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow GL(2,{\mathbb{F}})$ be a two-dimensional absolutely irreducible, continuous, odd representation of $G_{\mathbb{Q}}$. Here $\mathbb{F}$ is a finite field of characteristic $p$. Then $\overline{\rho}$ is modular, i.e., arises from (with respect to some fixed embedding $\imath:\overline{\mathbb{Q}}\hookrightarrow\overline{\mathbb{Q}_{p}})$ a newform $f$ of some weight $k\geq 2$ and level $N$ prime to $p$.
In 2009, Khare and Wintenberger [11, 12] proved that Serre’s Modularity Conjecture is true.
2. The Modularity of an Abelian Variety
Let $G:=Sp(2g,\mathbb{R})$ and $K=U(g).$ Let
$\mathbb{H}_{g}:=\{\,\Omega\in\mathbb{C}^{(g,g)}\,|\ \Omega=\,{}^{t}\Omega,\ \,{\rm{Im}}\,\Omega>0\ \}$
be the Siegel upper half plane of degree $g$. Then $G$ acts on ${\mathbb{H}}_{g}$ transitively by
$\alpha\cdot\Omega=(A\Omega+B)(C\Omega+D)^{-1},$
where $\alpha=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in G$ and $\Omega\in\mathbb{H}_{g}.$ The stabilizer of the action (2.1) at $iI_{g}$ is
$\left\{\begin{pmatrix}\,A&B\\ -B&A\end{pmatrix}\Big{|}\ A+iB\in U(g)\,\right\}\cong U(g).$
Thus we get the biholomorphic map
$G/K\longrightarrow\mathbb{H}_{g},\qquad\alpha K\mapsto\alpha\!\cdot\!iI_{g},\quad\alpha\in G.$
It is known that $\mathbb{H}_{g}$ is an Einstein-Kähler Hermitian symmetric space.
Let $\Gamma_{g}:=Sp(2g,\mathbb{Z})$ be the Siegel modular group of degree $g$. For a positive integer $N$, we let
$\Gamma_{g}(N):=\left\{\gamma\in\Gamma_{g}\,|\ \gamma\equiv I_{2g}\ ({\rm{mod}}\,N)\,\right\}$
be the the principal congruence subgroup of $\Gamma_{g}$ of level $N$. Let
$\Gamma_{g,0}(N):=\left\{\gamma\in\Gamma_{g}\,\big{|}\ \gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix},\quad C\equiv 0\ ({\rm{mod}}\,N)\,\right\}$
and
$\Gamma_{g,1}(N):=\left\{\gamma\in\Gamma_{g}\,\big{|}\ \gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\equiv\begin{pmatrix}I_{g}&*\\ 0&I_{g}\end{pmatrix}\ ({\rm{mod}}\,N)\,\right\}$
be the congruence subgroups of Level $N$. Then we have the relation
$\Gamma_{g}(N)\subset\Gamma_{g,1}(N)\subset\Gamma_{g,0}(N)\subset\Gamma_{g}.$
Definition 2.1.
Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$ and let $k$ be a nonnegative integer $k$. A function $F:\mathbb{H}_{g}\longrightarrow\mathbb{C}$ is called a $\sf{Siegel\ modular\ form}$ of degree $g$ and weight $k$ with respect to $\Gamma$ if it satisfies the following conditions : (S1) $F(\Omega)$ is holomorphic on $\mathbb{H}_{g}$ ; (S2) $F(\gamma\cdot\Omega)=(C\Omega+D)^{k}F(\Omega)$   for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma$ and $\Omega\in\mathbb{H}_{g}$ ; (S3) $F(\Omega)$ is bounded in any domain $Y\geq Y_{0}>0$ in the case $g=1$.
We denote the space of all Sigel modular forms of degree $g$ and weight $k$ with respect to $\Gamma$ by $[\Gamma,k]$. We define the so-called $\sf{Siegel\ operator}$
$\Phi_{g}:[\Gamma_{g},k]\longrightarrow[\Gamma_{g-1},k]$
by
$(\Phi_{g}(F))(\Omega_{1}):=\lim_{t\longrightarrow\infty}F\begin{pmatrix}\Omega_{1}&0\\ 0&i\,t\end{pmatrix},\quad\Omega_{1}\in\mathbb{H}_{g-1}.$
Then $\Phi_{g}$ is a well-defined linear mapping (cf, [13, pp. 187–189]). A Siegel modular form $F\in[\Gamma_{g},k]$ is called a $\sf{Siegel\ cusp\ form}$ if $\Phi_{g}(F)=0$ (cf. [13, p. 198]). We denote by $[\Gamma_{g},k]_{0}$ the space of all Siegel cusp forms in $[\Gamma_{g},k]$.
Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$. If $F\in[\Gamma,k]$, then $F$ has a Fourier expansion
$F(\Omega)=\sum_{T}a(T;F)\,e^{2\pi i\,{\rm Tr}(T\Omega)},$
where $T$ runs through all $g\times g$ half-integral semi-positive symmetric matrices. Here ${\rm Tr}(M)$ denotes the trace of a square matrix $M$. Following the Hecke’s method, Maass [13, pp. 202–210] associated with $F(\Omega)$ the Dirichlet series
$D(F,s):=\sum_{\{T\}}\frac{a(T;F)}{\varepsilon(T)}\,(\det T)^{-s},$
where the summation indicates that $T$ runs through a complete set of representatives of the sets
$\left\{\,T[U]\,|\ U\ \rm{unimodular}\,\right\},\ T>0$
and $\varepsilon(T)$ denotes the number of unimodular matrices $U$ which satisfy the equation $T[U]=T.$ We note that the numbers $\varepsilon(T)$ are finite.
Definition 2.2.
Let $F$ be a nonzero Siegel Hecke eigenform in $[\Gamma_{g},k]_{0}$. Let $\alpha_{p,0},\alpha_{p,1},\cdots,\alpha_{p,g}$ be the $p$-Satake parameters of $F$ at a prime $p$. We define the $\sf{local\ spinor\ zeta\ function}$ $Z_{F,p}(t)$ of $F$ at $p$ by
$Z_{F,p}(t):=(1-\alpha_{p,0}\,t)\sum_{r=1}^{g}\sum_{1\leq i_{1}<\cdots<i_{r}\leq g}(1-\alpha_{p,0}\alpha_{p,i_{1}}\cdots\alpha_{p,i_{r}}\,t).$
The $\sf{spinor\ zeta\ function}$ $Z_{F}(s)$ of $F$ is defined to be the following function
$Z_{F}(s):=\sum_{p:{\rm prime}}Z_{F,p}(p^{-s})^{-1},\quad{\rm Re}\,s\gg 0.$
Secondly one has the so-called $\sf{standard\ zeta\ function}$ $D_{F}(s)$ of a Siegel Hecke eigenform $F$ in $[\Gamma_{g},k]_{0}$ defined by
$D_{F}(s):=\sum_{p:{\rm prime}}D_{F,p}(p^{-s})^{-1},\quad{\rm Re}\,s\gg 0,$
where
$D_{F,p}(t)=(1-t)\sum_{i=1}^{g}(1-\alpha_{p,i}\,t)(1-\alpha_{p,i}^{-1}\,t).$
We refer to [23, p. 249].
Remark 2.3.
(1) If $g=1$, the spinor zeta function $Z_{f}(s)$ of a Hecke eigenform $f$ in $S_{k}(\Gamma_{1})$ is nothing but the Hecke $L$-function $L(f,s)$ of $f$. (2) If $g=1$, the standard zeta function $D_{f}(s)$ of a Hecke eigenform $f(\tau)=\sum_{n=1}^{\infty}a(n)\,e^{2\pi in\tau}$ in $S_{k}(\Gamma_{1})$ has the following equation
$D_{f}(s-k+1)=\sum_{p:{\rm prime}}(1+p^{k-s-1})^{-1}\cdot\sum_{n=1}^{\infty}a(n^{2})\,n^{-s}.$
Let $A$ be a $g$-dimensional simple abelian variety defined over $\mathbb{Q}$. For a prime $\ell$, we set
$A[\ell^{n}]:=\{\,x\in A(\overline{\mathbb{Q}})\,|\ \ell^{n}\cdot x=0\,\}.$
Then $A[\ell^{n}]\cong(\mathbb{Z}/\ell^{n}\mathbb{Z})^{g}\times(\mathbb{Z}/\ell^{n}\mathbb{Z})^{g}$ (cf. [15]). Then the Tate module of $A$ is given by
$T_{\ell}(A):=\lim_{\longleftarrow}A[\ell^{n}]\cong\mathbb{Z}_{\ell}^{g}\times\mathbb{Z}_{\ell}^{g}\cong\mathbb{Z}_{\ell}^{2g}.$
Therefore we have the $2g$-dimensional $\ell$-adic Galois representation of $G_{\mathbb{Q}}$
$\rho_{A,\ell}:G_{\mathbb{Q}}\longrightarrow GL(2g,\mathbb{Z}_{\ell})\subset GL(2g,\mathbb{Q}_{\ell}).$
Definition 2.4.
A $2g$-dimensional $\ell$-adic Galois representation $\rho$ of $G_{\mathbb{Q}}$ given by
$\rho:G_{\mathbb{Q}}\longrightarrow GL(2g,\mathbb{Z}_{\ell})\subset GL(2g,\mathbb{Q}_{\ell})$
is called $\sf{modular}$ if there is a Siegel Hecke eigenform $F(\Omega)\in[\Gamma_{g,0}(N),g+1]_{0}$ of weight $g+1$ with respect to $\Gamma_{g,0}(N)$ such that
${\rm Tr}\left(\rho({\rm Frob}_{p})\right)=a(pI_{g};F)\quad{\rm and}\quad\det\left(\rho({\rm Frob}_{p})\right)=p^{g}\quad{\rm for\ any\ prime}\ p\not|\,\ell N,$
where
$F(\Omega)=\sum_{T}a(T;F)\,e^{2\pi i\,{\rm Tr}(T\Omega)}$
is a Fourier expansion of $F(\Omega)$.
Definition 2.5.
Let $A$ be a $g$-dimensional simple abelian variety defined over $\mathbb{Q}$ and let $\ell$ be a prime. For a prime $p$, we let
$L_{p}(A,s):=\left\{\det\left(I_{2g}-p^{-s}\cdot\rho_{A,\ell}({\rm Frob}_{p})\big{|}_{T_{\ell}(A)}\right)\right\}^{-1}$
be the local $L$-function of $A$ at $p$. We define the $L$-function $L(A,s)$ of $A$ by
$L(A,s)=\prod_{p:\,{\rm prime}}L_{p}(A,s).$
Definition 2.6.
Let $A$ be a $g$-dimensional simple abelian variety defined over $\mathbb{Q}$. $A$ is called $\sf{modular}$ if there exists a Siegel Hecke eigenform $F(\Omega)\in[\Gamma_{g,0}(N),g+1]_{0}$ of weight $g+1$ with respect to $\Gamma_{g,0}(N)$ such that
$L(A,s)=D(F,s),\ Z_{F}(s)\ {\rm or}\ D_{F}(s).$
For two positive integers $g$ and $N$, we let
${\mathcal{A}}_{g,0}(N):=\Gamma_{g,0}(N)\backslash\mathbb{H}_{g}$
be the Siegel modular variety of level structure $N$ and let ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ be a smooth toroidal compactification of ${\mathcal{A}}_{g,0}(N)$ (cf. [1, 8]). We denote by
$\Omega^{i}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\quad 0\leq i\leq\frac{g(g+1)}{2}$
the complex vector space of holomorphic $i$-forms on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$. The Jacobian variety ${\rm Jac}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right)$ of ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ is defined to be the abelian variety
(2.1) ${{\rm Jac}}({\mathcal{A}}_{g,0}^{\rm tor}(N)):=\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right)^{\vee}/H_{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N),\mathbb{Z}\right),\qquad\nu=\frac{g(g+1)}{2}.$
The geometric genus of ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ is the dimension of the Jacobian variety ${\rm Jac}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right)$. It is known that the following two vector spaces are isomprphic :
(2.2) $[\Gamma_{g,0}(N),g+1]_{0}\cong\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad\nu=\frac{g(g+1)}{2}.$
More precisely, for a coordinate $\Omega=(\omega_{ij})\in\mathbb{H}_{g}$, we let
$\omega_{0}:=d\omega_{11}\wedge d\omega_{12}\wedge d\omega_{13}\wedge\cdots\wedge d\omega_{gg}$
be a holomorphic $\nu$-form on $\mathbb{H}_{g}$. If $\omega=F(\Omega)\,\omega_{0}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$, then
$F(\gamma\cdot\Omega)=\det(C\Omega+D)^{g+1}F(\Omega)$
for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{g,0}(N)$ and $\Omega\in\mathbb{H}_{g}.$ Thus $F\in[\Gamma_{g,0}(N),g+1].$ It was shown by Freitag [10] that $\omega$ can be extended to a holomorphic $\nu$-form on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ if and only if $F$ is a cusp form in $[\Gamma_{g,0}(N),g+1]_{0}.$ Indeed, the mapping
$[\Gamma_{g,0}(N),g+1]_{0}\longrightarrow\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad F\mapsto F\,\omega_{0}$
is an isomorphism as complex vector spaces. We observe that if $\omega_{k}:=G(\Omega)\,\omega_{0}^{\otimes k}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$ of degree $k\nu$, then $G(\Omega)\in[\Gamma_{g,0}(N),k(g+1)]_{0}$ is a cusp form of weight $k(g+1).$ Therefore according to (2.1) and (2.2), we have
(2.3) ${{\rm Jac}}({\mathcal{A}}_{g,0}^{\rm tor}(N))\cong[\Gamma_{g,0}(N),g+1]_{0}^{\vee}/H_{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N),\mathbb{Z}\right),\qquad\nu=\frac{g(g+1)}{2}.$
If there is no confusion, we simply set
$J_{g,0}(N):={{\rm Jac}}({\mathcal{A}}_{g,0}^{\rm tor}(N)).$
We propose the following conjectures.
Conjecture 2.7.
A simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$ is modular.
Conjecture 2.8.
Let A be a simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$. The following statements are equivalent : (MAV1) $A$ is modular. (MAV2) There exists a non-constant holomorphic mapping ${\mathcal{A}}_{g,0}^{\rm tor}(N)\longrightarrow A$ for some
positive integer $N$. (MAV3) There exists a non-constant holomorphic mapping $J_{g,0}(N)\longrightarrow A$ for some
positive integer $N$. (MAV4) $\rho_{A,\ell}$ is modular for any prime $\ell$.
We propose the following problems.
Problem 2.9.
Let $F\in[\Gamma_{g,0}(N),g+1]_{0}$ be a Siegel Hecke eigenform of weight $g+1$. Associate to $F$ a $2g$-dimensional continuous irreducible Galois representation of $G_{\mathbb{Q}}$.
Problem 2.10.
Let $k$ be a positive integer. Let $F\in[\Gamma_{g,0}(N),k]_{0}$ be a Siegel Hecke eigenform of weight $k$. Associate to $F$ a $2g$-dimensional continuous irreducible Galois representation of $G_{\mathbb{Q}}$.
Remark 2.11.
As mentioned in Section 1, in the case $g=1$, to a Hecke eigenform of weight $k\geq 1$, Shimura, Deligne and Serre associated a two-dimensional continuous irreducible Galois representation. For the case $g=2$, Taylor [19, 20] tried to associate the four dimensional continuous Galois representation of $G_{\mathbb{Q}}$ to a Siegel Hecke eigenform of small weight. But he did not specify the precise weight.

### v2

@@
 We conjecture that a simple abelian variety over $\mathbb{Q}$ is modular.
-The detailed article will appear later.
-Subject Classification: Primary 14Kxx, 14G35, 11F46, 11F80
-Keywords and phrases: modularity, abelian varieties, Siegel modular variety, Siegel modular forms, Galois
-representation, zeta functions
+Subject Classification: Primary 14Kxx, 14G35, 11F46, 11F40, 11F80
+Keywords and phrases: modularity, abelian variety, Siegel modular variety, Galois representation,
+Siegel Hecke eigenforms, zeta functions, Hecke algebra, $p$-Satake parameters.
+1. Introduction
+An elliptic curve $E$ over $\mathbb{Q}$ is said to be $\sf modular$} if the $L$-function $L(E,s)$ of $E$ equals the $L$-function $L(f,s)$ for some eigenform $f$, equivalently if $E$ has a finite covering by a modular curve of the form $X_{0}(N)$.
+At the Tokyo-Nikko conference held in 1955, Yutaka Taniyama (1927–1958) made a suggestion that every elliptic curve over $\mathbb{Q}$ is modular.
+At that time his suggestion was not clear and hence was not known in the mathematics community.
+In the early 1960’s, Goro Shimura (1930–2019) refined Taniyama’s suggestion through private conversations with a number of mathematicians and their efforts.
+In particular he discussed this subject with André Weil (1906–1998) seriously and intensively.
+Weil gave conceptual evidence for Taniyama’s suggestion in his famous paper [1] published in 1967.
+Through Weil’s paper, this suggestion was widely known as the so-called Shimura-Taniyama conjecture in the mathematics community.
+The Shimura-Taniyama conjecture associates objects of representation theory to those of algebraic geometry.
+It states that the $L$-series of an elliptic curve over $\mathbb{Q}$ which measures the behaviour of the curve mod $p$ for all primes $p$, can be identified with an integral transform of the Fourier series defined from an eigenform.
+In 1985 Frey [2] made the remarkable observation that the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ should imply Fermat’s Last Theorem.
+The precise mechanism relating the two was formulated by Serre [3] as the $\epsilon$-conjecture and this conjecture was proved by Ribet [4] in the summer of 1986.
+Ribet’s result only requires to prove the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ in order to reduce Fermat’s Last Theorem.
+As soon as Andrew Wiles learned Ribet’s result, he began to work the Shimura-Taniyama conjecture for a semistable elliptic curve over $\mathbb{Q}$ in the late summer of 1986.
+Here the semistabilty of an elliptic curve $E$ over $\mathbb{Q}$ is defined as follows :
+Definition 1.1.
+An elliptic curve $E$ over $\mathbb{Q}$ is said to be semistable at the prime $q$ if it is isomorphic to an elliptic curve $\tilde{E}$ over $\mathbb{Q}$ which $\tilde{E}$(mod $q$) is either nonsingular or has a node.
+An elliptic curve over $\mathbb{Q}$ is called semistable if it is semistable at every prime.
+On June 21-23 in 1993, Wiles had given a series of lectures under the title, ”Elliptic curves, modular forms and Galois representations” at the Issac Newton Institute for Mathematical Sciences in Cambridge, England.
+In this last lecture on June 23, Wiles commented that he had proved a part of the Shimura-Taniyama conjecture to the effect that every semistable elliptic curve over $\mathbb{Q}$ is modular.
+With the aide of the works of Frey [2], Serre [3] and Ribet [4], his proof solves Fermat’s Last Theorem which had been unsolved for more than 350 years.
+The news spread out all over the world through the well-known newspapers and magazines because Fermat’s Last Theorem holds great fascination for amateurs and professionals alike.
+But in the fall that year it turned out that the proof of Wiles was incomplete and flawed.
+Precisely, his construction of the Euler system used to extend Flach’s method was not complete.
+He was very concerned with filling the gaps in his flawed proof.
+At that time he was completely isolated from outside.
+In January 1994, he proposed Dr.
+Richard Taylor, his former Ph.D.
+student in Cambridge University, UK to join him in the attempt to repair the Euler system argument.
+Wiles still was convinced that his method was correct.
+Dr.
+Taylor accepted his proposal and joined in that project.
+On September 19th in 1994, Wiles was quite convinced that his method was correct and his gap could be filled up.
+After he invited Dr.
+Taylor to Princeton again, he completed his proof with the aid of Dr.
+Taylor on October, 1994 and submitted his paper to Annals of Mathematics on October 14, 1994.
+Finally his paper was accepted and was published in May, 1995 (cf.
+[5, 6]).
+We refer to [7] for more interesting stories.
+The aim of this article is to introduce the concept of the modularity of an abelian variety defined over the rational number field extending the modularity of an elliptic curve and to discuss the modularity of an abelian variety over $\mathbb{Q}$.
+An abelian variety is defined as follows in [8, p.
+39] :
+Definition 1.2.
+An abelian variety $X$ is a complete [^1] This means, in particular, that it is irreducible.
+See the footnote in [8, p.
+39].
+algebraic variety over an algebraically closed field $K$ with a group law $m_{X}:X\times X\longrightarrow X$ such that both $m_{X}$ and the inverse map are morphisms of varieties.
+An abelian variety over $\mathbb{Q}$ is a connected and complex projective manifold that is also a group variety generalizing elliptic curves over $\mathbb{Q}$.
+The paper is organized as follows.
+In Section 2, we briefly outline the modularity of an elliptic curve over $\mathbb{Q}$.
+Wiles proved the modularity of a semistable elliptic curve over $\mathbb{Q}$ leading to solve Fermat’s Last Theorem (cf.
+[5, 6]).
+In 2001, Breuil, Conrad, Diamond and Taylor proved that every elliptic curve over $\mathbb{Q}$ is modular (cf.
+[9]).
+In Section 3, we introduce the notion of the modularity of an abelian variety over $\mathbb{Q}$.
+We propose several conjectures and open problems.
+In Section 4 (Appendix), we describe and survey the Hecke algebra for a symplectic group and $p$-Satake parameters for the readers.
+Notations: We denote by $\mathbb{Q},\,\mathbb{R}$ and $\mathbb{C}$ the field of rational numbers, the field of real numbers and the field of complex numbers respectively.
+We denote by $\mathbb{Z}$ and $\mathbb{Z}^{+}$ the ring of integers and the set of all positive integers respectively.
+For a prime $p$, we denote by $\mathbb{Q}_{p}$ the field of $p$-adic numbers and by $\mathbb{Z}_{p}$ the ring of $p$-adic integers.
+$\overline{\mathbb{Q}}$ (resp.
+$\overline{\mathbb{Q}_{p}}$) denotes the algebraic closure of $\mathbb{Q}$ (resp.
+$\mathbb{Q}_{p}$).
+$G_{\mathbb{Q}}={\rm Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denotes the absolute Galois group of $\mathbb{Q}$.
+The symbol “:=” means that the expression on the right is the definition of that on the left.
+For two positive integers $k$ and $l$, $F^{(k,l)}$ denotes the set of all $k\times l$ matrices with entries in a commutative ring $F$.
+For a square matrix $M\in F^{(k,k)}$ of degree $k$, ${\rm Tr}(M)$} denotes the trace of $M$.
+For any $M\in F^{(k,l)},\ ^{t}\!M$ denotes the transpose of a matrix $M$.
+$I_{n}$ denotes the identity matrix of degree $n$.
+For $A\in F^{(k,l)}$ and $B\in F^{(k,k)}$, we set $B[A]=\,^{t}ABA$ (Siegel’s notation).
+For a number field $F$, we denote by ${\mathbb{A}}_{F}$ the ring of adeles of $F$.
+If $F=\mathbb{Q}$, the subscript will be omitted.
+$Sp(2g,\mathbb{R})=\left\{\,M\in Sp(2g,\mathbb{R})\,|\ {}^{t}MJ_{g}M=J_{g}[M]=J_{g}\,\right\}$
+denotes the symplectic group of degree $g$, where
+$J_{g}=\begin{pmatrix}0&I_{g}\\ -I_{g}&0\end{pmatrix}$.
+$\Gamma_{g}:=Sp(2g,\mathbb{Z})=\left\{\ \begin{pmatrix}A&B\\ C&D\end{pmatrix}\in Sp(2g,\mathbb{R})\Big|\ A,B,C,D\ \rm{integral}\,\right\}$
+denotes the Siegel modular group of degree $g$.
+For a prime $\ell$ and a positive integer $n$, ${\mathbb{F}}_{\ell^{n}}$ denotes the field with $\ell^{n}$ elements.
+The algebraic closure $\overline{\mathbb{F}}_{\ell}$ of $\mathbb{F}$$_{\ell}$ is given by
+$\overline{\mathbb{F}}_{\ell}=\bigcup_{n\geq 1}{\mathbb{F}}_{\ell^{n}}.$
+If $p$ is prime and $\lambda$ is a prime ideal dividing $p$ in the ring of integers in $\overline{\mathbb{Q}}$, there exist a filtration
+$I_{\lambda}\subset D_{\lambda}\subset G_{\mathbb{Q}},$
+where the decomposition group $D_{\lambda}$ and the inertia group $I_{\lambda}$ are defined respectively by
+$D_{\lambda}:=\{\,\sigma\in G_{\mathbb{Q}}\,|\,\sigma(\lambda)=\lambda\,\}$
+and
+$I_{\lambda}:=\{\,\sigma\in D_{\lambda}\,|\,\sigma(x)\equiv x\,({\rm mod}\,\lambda)\ {\rm{for\ all\ algebraic\ integers}}\ x\,\}.$
+Then there are natural identifications
+$D_{\lambda}\cong{\rm Gal}(\overline{\mathbb{Q}}_{p}/\mathbb{Q}_{p})\quad{\rm and}\quad D_{\lambda}/I_{\lambda}\cong{\rm Gal}(\overline{\mathbb{F}}_{p}/\mathbb{F}_{p}).$
+We denote by ${\rm Frob}_{\lambda}\in D_{\lambda}/I_{\lambda}$ the inverse image of the canonical generator $x\mapsto x^{p}$ of ${\rm Gal}(\overline{\mathbb{F}}_{p}/{\mathbb{F}}_{p}).$ If $\lambda^{\prime}$ is another prime lying above $p$, then $\lambda^{\prime}=\sigma(\lambda)$ for some $\sigma\in G_{\mathbb{Q}}$ and
+$D_{\lambda^{\prime}}=\sigma D_{\lambda}\sigma^{-1},\quad I_{\lambda^{\prime}}=\sigma I_{\lambda}\sigma^{-1}\quad{\rm and}\quad{\rm Frob}_{\lambda^{\prime}}=\sigma\,{\rm Frob}_{\lambda}\,\sigma^{-1}.$
+Since we will care about these objects only up to conjugation, we will write $D_{\lambda}$ and $I_{\lambda}$.
+Now we will write ${\rm Frob}_{p}$ for any representative of a ${\rm Frob}_{\lambda}$.
+If $\rho$ is a representation of $G_{\mathbb{Q}}$ which is unramified at $p$, then ${\rm Tr}(\rho({\rm Frob}_{p}))$ and $\det(\rho({\rm Frob}_{p}))$ are well defined, that is, are independent of the choice of $\lambda$.
 2. The Modularity of an Elliptic Curve
 We set $\Gamma_{1}:=SL(2,\mathbb{Z})$.
-For a positive integer $N$, we let $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$ be the congruence subgroups of $\Gamma_{1}$ such that $\Gamma(N)\subset\Gamma_{1}(N)\subset\Gamma_{0}(N)\subset\Gamma_{1}.$ We refer to [6, pp.
+For a positive integer $N$, we let $\Gamma(N),\ \Gamma_{1}(N)$ and $\Gamma_{0}(N)$ be the congruence subgroups of $\Gamma_{1}$ such that $\Gamma(N)\subset\Gamma_{1}(N)\subset\Gamma_{0}(N)\subset\Gamma_{1}.$ We refer to [10, pp.
 13–14, p.
@@
 $Y_{0}(N)$) as a dense Zariski open subset (cf.
-see [6, pp.
+see [10, pp.
 45–60]).
-Let $S_{k}(N)$ be the space of cusp forms of weight $k\geq 1$ and level $N\geq 1$.
-Here $k$ and $N$ be positive integers.
-We recall that if $f\in S_{k}(N)$, it satisfies the following properties : (C1) $f((a\tau+b)(c\tau+d)^{-1})=(c\tau+d)^{k}f(\tau)$ for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}(N)$ and $\tau\in\mathbb{H}_{1}$; (C2) $|f(\tau)|^{2}({\rm Im}\,\tau)^{k}$ is bounded in $\mathbb{H}_{1}$ and (C3) the Fourier expansion of $f(\tau)$ is given by
+Let $S_{k}(N)$ be the space of cusp forms of weight $k\geq 1$ and level $N\geq 1$. Here $k$ and $N$ be positive integers. We recall that if $f\in S_{k}(N)$, it satisfies the following properties :
+(C1) $f((a\tau+b)(c\tau+d)^{-1})=(c\tau+d)^{k}f(\tau)$ for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}(N)$ and $\tau\in\mathbb{H}_{1}$;
+(C2) $|f(\tau)|^{2}({\rm Im}\,\tau)^{k}$ is bounded in $\mathbb{H}_{1}$ and
+(C3) the Fourier expansion of $f(\tau)$ is given by
 $f(\tau)=\sum_{n=1}^{\infty}a_{n}(f)\,q^{n},\quad{\rm where}\ q=e^{2\pi i\tau}.$
@@
 $L(f,s):=\sum_{n=1}^{\infty}a_{n}(f)\,n^{-s}.$
-For each prime $p\!\not|N$, we recall that the Hecke operator $T_{p}:S_{k}(N)\longrightarrow S_{k}(N)$ is defined by
+[…]
 $(T_{p}f)(\tau)=p^{-1}\sum_{i=0}^{p-1}f\left(\frac{\tau+i}{p}\right)+p^{k-1}f\left(\frac{ap\tau+b}{cp\tau+d}\right),\quad f\in S_{k}(N)$
-for any $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{1}$ with $c\equiv 0\,({\rm mod}\,N)$ and $d\equiv p\,({\rm mod}\,N)$.
-We refer to [2, p.
-844] or [6, pp.
-170–171] for more detail.
-The Hecke operators $T_{p}\,(p\!\!\not|N)$ can be simultaneously diagonalized on $S_{k}(N)$ and a simultaneous eigenvector a Hecke eigenform or simply an eigenform.
-Let $\lambda$ be a place of the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ lying over a rational integer $\ell$ and ¯${\bar{\mathbb{Q}}}_{\lambda}$ denote the algebraic closure of $\mathbb{Q}_{\ell}$ via $\lambda$.
+[…]
+Let $\lambda$ be a place of the algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ in $\mathbb{C}$ lying over a rational integer $\ell$ and ${\bar{\mathbb{Q}}}_{\lambda}$ denote the algebraic closure of $\mathbb{Q}_{\ell}$ via $\lambda$.
 Let $G_{\mathbb{Q}}:={\rm Gal}({\bar{\mathbb{Q}}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$.
@@
 $\rho_{f,\lambda}:G_{\mathbb{Q}}\longrightarrow GL(2,{\bar{\mathbb{Q}}}_{\lambda})$
-such that $\rho_{f,\lambda}$ is unramified at $p$ for all primes $p\!\!\not|\,\ell N$ and
-${\rm Tr}\left(\rho_{f,\lambda}({\rm Frob}_{p})\right)=a_{p}(f)\quad{\rm for\ any\ prime}\ p\!\!\not|\,\ell N.$
+[…]
+[…]
 The existence of $\rho_{f,\lambda}$ is due to Shimura if $k=2$ [11], due to Deligne if $k>2$ [12] and due to Deligne and Serre if $k=1$ [13].
@@
 The $L$-function $L(E,s)$ of $E$ is defined by the product of the local $L$-factors
-$L(E,s):=\prod_{p|D}\left(\frac{1}{1-a_{p}(E)p^{-s}}\right)\prod_{p\not|D}\left(\frac{1}{1-a_{p}(E)\,p^{-s}+p^{1-2s}}\right).$
+[…]
 Then $L(E,s)$ converges absolutely for ${\rm Re}\,s>\frac{3}{2}$ and extends to an entire function by [9].
@@
 Here $\Omega^{1}(X_{1}(N))$ denotes the complex vector space of holomorphic 1-forms on $X_{1}(N)$ and $W^{\vee}$ denotes the dual space of a complex vector space $W$.
-It is known that the following statements are equivalent : (a) $E$ is modular.
-(b) There is a non-constant holomorphic mapping $X_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
-integer $N$. (c) There is a non-constant holomorphic mapping $J_{1}(N)\longrightarrow E(\mathbb{C})$ for some positive
-integer $N$. (d) $\rho_{E,\ell}$ is modular for a prime $\ell$.
+It is known that the following statements are equivalent :
+(a) $E$ is modular.
+(b) There is a non-constant holomorphic mapping $X_{1}(N)\longrightarrow E(\mathbb{C})$ for some pos-
+itive integer $N$.
+(c) There is a non-constant holomorphic mapping $J_{1}(N)\longrightarrow E(\mathbb{C})$ for some pos-
+itive integer $N$.
+(d) $\rho_{E,\ell}$ is modular for a prime $\ell$.
 The above statements have been called the Taniyama-Shimura conjecture.
@@
 The implication (c) $\Longrightarrow$ (b) is obvious.
-Wiles [21, 22] proved that a semistable eliptic curve over $\mathbb{Q}$ is modular by proving the statement (d).
-Thereafter Breuil, Conrad, Diamond and Taylor [2] proved that every eliptic curve over $\mathbb{Q}$ is modular.
-Serre [16] conjectured the following : Serre’s Modularity Conjecture : Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow GL(2,{\mathbb{F}})$ be a two-dimensional absolutely irreducible, continuous, odd representation of $G_{\mathbb{Q}}$.
+Wiles [5, 6] proved that a semistable elliptic curve over $\mathbb{Q}$ is modular by proving the statement (d).
+Thereafter Breuil, Conrad, Diamond and Taylor [9] proved that every elliptic curve over $\mathbb{Q}$ is modular.
+Serre [3] conjectured the following :
+Serre’s Modularity Conjecture : Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow GL(2,{\mathbb{F}})$ be a two-dimensional absolutely irreducible, continuous, odd representation of $G_{\mathbb{Q}}$.
 Here $\mathbb{F}$ is a finite field of characteristic $p$.
@@
 be the Siegel upper half plane of degree $g$. Then $G$ acts on ${\mathbb{H}}_{g}$ transitively by
-$\alpha\cdot\Omega=(A\Omega+B)(C\Omega+D)^{-1},$
+(3.1) $\alpha\cdot\Omega=(A\Omega+B)(C\Omega+D)^{-1},$
 where $\alpha=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in G$ and $\Omega\in\mathbb{H}_{g}.$ The stabilizer of the action (3.1) at $iI_{g}$ is
@@
 Thus we get the biholomorphic map
-$G/K\longrightarrow\mathbb{H}_{g},\qquad\alpha K\mapsto\alpha\!\cdot\!iI_{g},\quad\alpha\in G.$
+$G/K\longrightarrow\mathbb{H}_{g},\qquad\alpha K\longmapsto\alpha\!\cdot\!iI_{g},\quad\alpha\in G.$
 It is known that $\mathbb{H}_{g}$ is an Einstein-Kähler Hermitian symmetric space.
@@
 Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$ and let $k$ be a nonnegative integer $k$.
-A function $F:\mathbb{H}_{g}\longrightarrow\mathbb{C}$ is called a $\sf{Siegel\ modular\ form}$ of degree $g$ and weight $k$ with respect to $\Gamma$ if it satisfies the following conditions : (S1) $F(\Omega)$ is holomorphic on $\mathbb{H}_{g}$ ; (S2) $F(\gamma\cdot\Omega)=(C\Omega+D)^{k}F(\Omega)$   for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma$ and $\Omega\in\mathbb{H}_{g}$ ; (S3) $F(\Omega)$ is bounded in any domain $Y\geq Y_{0}>0$ in the case $g=1$.
-We denote the space of all Sigel modular forms of degree $g$ and weight $k$ with respect to $\Gamma$ by $[\Gamma,k]$. We define the so-called $\sf{Siegel\ operator}$
+A function $F:\mathbb{H}_{g}\longrightarrow\mathbb{C}$ is called a $\sf{Siegel\ modular\ form}$ of degree $g$ and weight $k$ with respect to $\Gamma$ if it satisfies the following conditions :
+(S1) $F(\Omega)$ is holomorphic on $\mathbb{H}_{g}$ ;
+(S2) $F(\gamma\cdot\Omega)=(C\Omega+D)^{k}F(\Omega)$ for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma$ and $\Omega\in\mathbb{H}_{g}$ ;
+(S3) $F(\Omega)$ is bounded in any domain $Y\geq Y_{0}>0$ in the case $g=1$.
+We denote the space of all Siegel modular forms of degree $g$ and weight $k$ with respect to $\Gamma$ by $\Gamma,k]$.
+We define the so-called $\sf{Siegel\ operator}$
 $\Phi_{g}:[\Gamma_{g},k]\longrightarrow[\Gamma_{g-1},k]$
@@
 $(\Phi_{g}(F))(\Omega_{1}):=\lim_{t\longrightarrow\infty}F\begin{pmatrix}\Omega_{1}&0\\ 0&i\,t\end{pmatrix},\quad\Omega_{1}\in\mathbb{H}_{g-1}.$
-Then $\Phi_{g}$ is a well-defined linear mapping (cf, [13, pp.
+Then $\Phi_{g}$ is a well-defined linear mapping (cf, [21, pp.
 187–189]).
 A Siegel modular form $F\in[\Gamma_{g},k]$ is called a $\sf{Siegel\ cusp\ form}$ if $\Phi_{g}(F)=0$ (cf.
-[13, p.
+[21, p.
 198]).
-We denote by $[\Gamma_{g},k]_{0}$ the space of all Siegel cusp forms in $[\Gamma_{g},k]$.
+We denote by $\Gamma_{g},k]_{0}$ the space of all Siegel cusp forms in $\Gamma_{g},k]$.
 Let $\Gamma$ be a congruence subgroup of $\Gamma_{g}$. If $F\in[\Gamma,k]$, then $F$ has a Fourier expansion
@@
 Here ${\rm Tr}(M)$ denotes the trace of a square matrix $M$.
-Following the Hecke’s method, Maass [13, pp.
+Following Hecke’s method, Maass [21, pp.
 202–210] associated with $F(\Omega)$ the Dirichlet series
@@
 Definition 3.2.
-Let $F$ be a nonzero Siegel Hecke eigenform in $[\Gamma_{g},k]_{0}$.
-Let $\alpha_{p,0},\alpha_{p,1},\cdots,\alpha_{p,g}$ be the $p$-Satake parameters of $F$ at a prime $p$.
-We define the $\sf{local\ spinor\ zeta\ function}$ $Z_{F,p}(t)$ of $F$ at $p$ by
+Let $F$ be a nonzero Siegel Hecke eigenform in $\Gamma_{g},k]_{0}$. Let $\alpha_{p,0},\alpha_{p,1},\cdots,$
+$\alpha_{p,g}$ be the $p$-Satake parameters of $F$ at a prime $p$ (cf. see Section 4). We define the $\sf{local\ spinor\ zeta\ function}$ $Z_{F,p}(t)$ of $F$ at $p$ by
 $Z_{F,p}(t):=(1-\alpha_{p,0}\,t)\sum_{r=1}^{g}\sum_{1\leq i_{1}<\cdots<i_{r}\leq g}(1-\alpha_{p,0}\alpha_{p,i_{1}}\cdots\alpha_{p,i_{r}}\,t).$
@@
 $Z_{F}(s):=\sum_{p:{\rm prime}}Z_{F,p}(p^{-s})^{-1},\quad{\rm Re}\,s\gg 0.$
-Secondly one has the so-called $\sf{standard\ zeta\ function}$ $D_{F}(s)$ of a Siegel Hecke eigenform $F$ in $[\Gamma_{g},k]_{0}$ defined by
+Secondly one has the so-called $\sf{standard\ zeta\ function}$ $D_{F}(s)$ of a Siegel Hecke eigenform $F$ in $\Gamma_{g},k]_{0}$ defined by
 $D_{F}(s):=\sum_{p:{\rm prime}}D_{F,p}(p^{-s})^{-1},\quad{\rm Re}\,s\gg 0,$
@@
 $D_{F,p}(t)=(1-t)\sum_{i=1}^{g}(1-\alpha_{p,i}\,t)(1-\alpha_{p,i}^{-1}\,t).$
-We refer to [23, p. 249].
+We refer to [22, p. 249].
 Remark 3.3.
@@
 is called $\sf{modular}$ if there is a Siegel Hecke eigenform $F(\Omega)\in[\Gamma_{g,0}(N),g+1]_{0}$ of weight $g+1$ with respect to $\Gamma_{g,0}(N)$ such that
-${\rm Tr}\left(\rho({\rm Frob}_{p})\right)=a(pI_{g};F)\quad{\rm and}\quad\det\left(\rho({\rm Frob}_{p})\right)=p^{g}\quad{\rm for\ any\ prime}\ p\not|\,\ell N,$
+[…]
 where
@@
 For two positive integers $g$ and $N$, we let
-${\mathcal{A}}_{g,0}(N):=\Gamma_{g,0}(N)\backslash\mathbb{H}_{g}$
+[…]
 be the Siegel modular variety of level structure $N$ and let ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ be a smooth toroidal compactification of ${\mathcal{A}}_{g,0}(N)$ (cf. [23, 24]). We denote by
@@
 It is known that the following two vector spaces are isomprphic :
-(2.2) $[\Gamma_{g,0}(N),g+1]_{0}\cong\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad\nu=\frac{g(g+1)}{2}.$
+(3.3) $\Gamma_{g,0}(N),g+1]_{0}\cong\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad\nu=\frac{g(g+1)}{2}.$
 More precisely, for a coordinate $\Omega=(\omega_{ij})\in\mathbb{H}_{g}$, we let
@@
 $F(\gamma\cdot\Omega)=\det(C\Omega+D)^{g+1}F(\Omega)$
-for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{g,0}(N)$ and $\Omega\in\mathbb{H}_{g}.$ Thus $F\in[\Gamma_{g,0}(N),g+1].$ It was shown by Freitag [10] that $\omega$ can be extended to a holomorphic $\nu$-form on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ if and only if $F$ is a cusp form in $[\Gamma_{g,0}(N),g+1]_{0}.$ Indeed, the mapping
-$[\Gamma_{g,0}(N),g+1]_{0}\longrightarrow\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad F\mapsto F\,\omega_{0}$
+for all $\gamma=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Gamma_{g,0}(N)$ and $\Omega\in\mathbb{H}_{g}.$ Thus $F\in[\Gamma_{g,0}(N),g+1].$ It was shown by Freitag [25] that $\omega$ can be extended to a holomorphic $\nu$-form on ${\mathcal{A}}_{g,0}^{\rm tor}(N)$ if and only if $F$ is a cusp form in $\Gamma_{g,0}(N),g+1]_{0}.$ Indeed, the mapping
+$\Gamma_{g,0}(N),g+1]_{0}\longrightarrow\Omega^{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N)\right),\qquad F\longmapsto F\,\omega_{0}$
 is an isomorphism as complex vector spaces.
-We observe that if $\omega_{k}:=G(\Omega)\,\omega_{0}^{\otimes k}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$ of degree $k\nu$, then $G(\Omega)\in[\Gamma_{g,0}(N),k(g+1)]_{0}$ is a cusp form of weight $k(g+1).$ Therefore according to (2.1) and (2.2), we have
+We observe that if $\omega_{k}:=G(\Omega)\,\omega_{0}^{\otimes k}$ is a $\Gamma_{g,0}(N)$-invariant holomorphic form on $\mathbb{H}_{g}$ of degree $k\nu$, then $G(\Omega)\in[\Gamma_{g,0}(N),k(g+1)]_{0}$ is a cusp form of weight $k(g+1).$
+Therefore according to (3.2) and (3.3), we have
 (3.4) ${{\rm Jac}}({\mathcal{A}}_{g,0}^{\rm tor}(N))\cong[\Gamma_{g,0}(N),g+1]_{0}^{\vee}/H_{\nu}\left({\mathcal{A}}_{g,0}^{\rm tor}(N),\mathbb{Z}\right),\qquad\nu=\frac{g(g+1)}{2}.$
@@
 Conjecture 3.8.
-Let A be a simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$.
-The following statements are equivalent : (MAV1) $A$ is modular.
+Let A be a simple abelian variety of dimesion $g$ defined over $\mathbb{Q}$. The following statements are equivalent :
+(MAV1) $A$ is modular.
 (MAV2) There exists a non-constant holomorphic mapping ${\mathcal{A}}_{g,0}^{\rm tor}(N)\longrightarrow A$ for some
-positive integer $N$. (MAV3) There exists a non-constant holomorphic mapping $J_{g,0}(N)\longrightarrow A$ for some
-positive integer $N$. (MAV4) $\rho_{A,\ell}$ is modular for any prime $\ell$.
+positive integer $N$.
+(MAV3) There exists a non-constant holomorphic mapping $J_{g,0}(N)\longrightarrow A$ for some
+positive integer $N$.
+(MAV4) $\rho_{A,\ell}$ is modular for any prime $\ell$.
 We propose the following problems.
-Problem 2.9.
+Problem 3.9.
 Let $F\in[\Gamma_{g,0}(N),g+1]_{0}$ be a Siegel Hecke eigenform of weight $g+1$. Associate to $F$ a $2g$-dimensional continuous irreducible Galois representation of $G_{\mathbb{Q}}$.
-Problem 2.10.
+Problem 3.10.
 Let $k$ be a positive integer.
@@
 Remark 3.11.
-As mentioned in Section 1, in the case $g=1$, to a Hecke eigenform of weight $k\geq 1$, Shimura, Deligne and Serre associated a two-dimensional continuous irreducible Galois representation.
+As mentioned in Section 2, in the case $g=1$, to a Hecke eigenform of weight $k\geq 1$, Shimura, Deligne and Serre [11, 12, 13] associated a two-dimensional continuous irreducible Galois representation.
 For the case $g=2$, Taylor [26, 27] tried to associate the four dimensional continuous Galois representation of $G_{\mathbb{Q}}$ to a Siegel Hecke eigenform of small weight.
 But he did not specify the precise weight.
+4. Appendix : The Hecke Algebra and $p$-Satake Parameters
+4.1. The Structure of the Hecke Algebra
+For a positive integer $g$, we let $\Gamma_{g}=Sp(2g,\mathbb{Z})$ and let
+$\Delta_{g}:=GSp(2g,\mathbb{Q})=\big\{\,M\in GL(2g,\mathbb{Q})\,|\ {}^{t}MJ_{g}M=l(M)J_{g},\ l(M)\in\mathbb{Q}^{\times}\,\big\}$
+be the group of symplectic similitudes of the rational symplectic vector space $(\mathbb{Q}^{2g},\langle\ ,\ \rangle)$. We put
+$\Delta_{g}^{+}:=GSp(2g,\mathbb{Q})^{+}=\big\{\,M\in\Delta_{g}\,|\ l(M)>0\,\big\}.$
+Following the notations in [25], we let ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ be the complex vector space of all formal finite sums of double cosets $\Gamma_{g}M\Gamma_{g}$ with $M\in\Delta_{g}^{+}$.
+A double coset $\Gamma_{g}M\Gamma_{g}\,(M\in\Delta_{g}^{+})$ can be written as a finite disjoint union of right cosets $\Gamma_{g}M_{\nu}\,(1\leq\nu\leq h)\,:$
+$\Gamma_{g}M\Gamma_{g}=\bigcup^{h}_{\nu=1}\Gamma_{g}M_{\nu}\quad(\textrm{disjoint}).$
+Let ${\mathscr{L}}(\Gamma_{g},\Delta_{g})$ be the complex vector space consisting of formal finite sums of right cosets $\Gamma_{g}M$ with $M\in\Delta^{+}$.
+For each double coset $\Gamma_{g}M\Gamma_{g}=\bigcup^{h}_{\nu=1}\Gamma_{g}M_{\nu}$ we associate an element $j(\Gamma_{g}M\Gamma_{g})$ in ${\mathscr{L}}(\Gamma_{g},\Delta_{g})$ defined by
+$j(\Gamma_{g}M\Gamma_{g}):=\sum^{h}_{\nu=1}\Gamma_{g}M_{\nu}.$
+Then $j$ induces a linear map
+(4.1) $j_{*}:{\mathscr{H}}(\Gamma_{g},\Delta_{g})\longrightarrow{\mathscr{L}}(\Gamma_{g},\Delta_{g}).$
+We observe that $\Delta_{g}$ acts on ${\mathscr{L}}(\Gamma_{g},\Delta_{g})$ as follows:
+$\big(\sum_{j=1}^{h}c_{j}\,\Gamma_{g}M_{j}\big)\cdot M=\sum_{j=1}^{h}c_{j}\,\Gamma_{g}M_{j}M,\quad M\in\Delta_{g}.$
+We denote
+${\mathscr{L}}(\Gamma_{g},\Delta_{g})^{\Gamma_{g}}:=\big\{\,T\in{\mathscr{L}}(\Gamma_{g},\Delta_{g})\,|\ T\cdot\gamma=T\ \textrm{for all}\ \gamma\in\Gamma_{g}\,\big\}$
+be the subspace of $\Gamma_{g}$-invariants in ${\mathscr{L}}(\Gamma_{g},\Delta_{g})$.
+Then we can show that ${\mathscr{L}}(\Gamma_{g},\Delta_{g})^{\Gamma_{g}}$ coincides with the image of $j_{*}$ and the map
+(4.2) $j_{*}:{\mathscr{H}}(\Gamma_{g},\Delta_{g})\longrightarrow{\mathscr{L}}(\Gamma_{g},\Delta_{g})^{\Gamma_{g}}$
+is an isomorphism of complex vector spaces (cf. [25, p. 228]). From now on we identify ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ with ${\mathscr{L}}(\Gamma_{g},\Delta_{g})^{\Gamma_{g}}.$
+We define the multiplication of the double coset $\Gamma_{g}M\Gamma_{g}$ and $\Gamma_{g}N$ by
+(4.3) $(\Gamma_{g}M\Gamma_{g})\cdot(\Gamma_{g}N)=\sum_{j=1}^{h}\Gamma_{g}M_{j}N,\quad M,N\in\Delta_{g},$
+where $\Gamma_{g}M\Gamma_{g}=\bigcup^{h}_{j=1}\Gamma_{g}M_{j}\ (\textrm{disjoint}).$ The definition (4.3) is well defined, i.e., independent of the choice of $M_{j}$ and $N$.
+We extend this multiplication to ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ and ${\mathscr{L}}(\Gamma_{g},\Delta_{g})$.
+Since
+${\mathscr{H}}(\Gamma_{g},\Delta_{g})\cdot{\mathscr{H}}(\Gamma_{g},\Delta_{g})\subset{\mathscr{H}}(\Gamma_{g},\Delta_{g}),$
+${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ is an associative algebra with the identity element $\Gamma_{g}I_{2g}\Gamma_{g}=\Gamma_{g}$.
+The algebra ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ is called the Hecke algebra with respect to $\Gamma_{g}$ and $\Delta_{g}$.
+We now describe the structure of the Hecke algebra ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$.
+For a prime $p$, we let $\mathbb{Z}[1/p]$ be the ring of all rational numbers of the form $a\cdot p^{\nu}$ with $a,\nu\in\mathbb{Z}.$ For a prime $p$, we denote
+$\Delta_{g,p}:=\Delta_{g}\cap GL\big(2g,\mathbb{Z}[1/p]\big).$
+Then we have a decomposition of ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$
+${\mathscr{H}}(\Gamma_{g},\Delta_{g})=\bigotimes_{p\,:\,\textrm{prime}}{\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$
+as a tensor product of local Hecke algebras ${\mathscr{H}}(\Gamma_{g},\Delta_{g,p}).$ We denote by $\check{\mathscr{H}}(\Gamma_{g},\Delta_{g})$ (resp.
+$\check{\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$ the subring of ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ (resp.
+${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$ by integral matrices.
+In order to describe the structure of local Hecke operators ${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$, we need the following lemmas.
+Lemma 4.1.
+Let $M\in\Delta_{g}^{+}$ with $^{t}MJ_{g}M=lJ_{g}$. Then the double coset $\Gamma_{g}M\Gamma_{g}$ has a unique representative of the form
+$M_{0}={\rm diag}(a_{1},\cdots,a_{g},d_{1},\cdots,d_{g}),$
+where $a_{g}|d_{g},\ a_{j}>0,\ a_{j}d_{j}=l$ for $1\leq j\leq g$ and $a_{k}|a_{k+1}$ for $1\leq k\leq g-1.$
+For a positive integer $l$, we let
+$O_{g}(l):=\big\{\,M\in GL(2g,\mathbb{Z})\,|\ {}^{t}MJ_{g}M=lJ_{g}\ \big\}.$
+Then we see that $O_{g}(l)$ can be written as a finite disjoint union of double cosets and hence as a finite union of right cosets.
+We define $T(l)$ as the element of ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ defined by $O_{g}(l).$
+Lemma 4.2.
+(a) Let $l$ be a positive integer. Let
+$O_{g}(l)=\cup_{\nu=1}^{h}\Gamma_{g}M_{\nu}\quad({\rm disjoint})$
+be a disjoint union of right cosets $\Gamma_{g}M_{\nu}\,(1\leq\nu\leq h).$ Then each right coset $\Gamma_{g}M_{\nu}$ has a representative of the form
+$M_{\nu}=\begin{pmatrix}A_{\nu}&B_{\nu}\\ 0&D_{\nu}\end{pmatrix},\quad{}^{t}A_{\nu}D_{\nu}=l\,I_{g},\quad A_{\nu}\ {\rm is\ upper\ triangular}.$
+(b) Let $p$ be a prime. Then
+$T(p)=O_{g}(p)=\Gamma_{g}\begin{pmatrix}I_{g}&0\\ 0&pI_{g}\end{pmatrix}\Gamma_{g}$
+and
+$T(p^{2})=\sum_{i=0}^{g}T_{i}(p^{2}),$
+where
+$T_{k}(p^{2}):=\begin{pmatrix}I_{g-k}&0&0&0\\ 0&pI_{k}&0&0\\ 0&0&p^{2}I_{g-k}&0\\ 0&0&0&pI_{k}\end{pmatrix}\Gamma_{g},\quad 0\leq k\leq g.$
+Proof. The proof can be found in [25, p. 225 and p. 250]. $\square$
+For example, $T_{g}(p^{2})=\Gamma_{g}(pI_{2g})\Gamma_{g}$ and
+$T_{0}(p^{2})=\Gamma_{g}\begin{pmatrix}I_{g}&0\\ 0&p^{2}I_{g}\end{pmatrix}\Gamma_{g}=T(p)^{2}.$
+We have the following
+Theorem 4.3.
+The local Hecke algebra $\check{\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$ is generated by algebraically independent generators $T(p),\,T_{1}(p^{2}),\cdots,T_{g}(p^{2}).$
+Proof. The proof can be found in [25, p. 250 and p. 261]. $\square$
+On $\Delta_{g}$ we have the anti-automorphism $M\mapsto M^{*}:=l(M)M^{-1}\,(M\in\Delta_{g})$.
+Obviously $\Gamma_{g}^{*}=\Gamma_{g}$.
+By Lemma 4.1, $(\Gamma_{g}M\Gamma_{g})^{*}=\Gamma_{g}M^{*}\Gamma_{g}=\Gamma_{g}M\Gamma_{g}.$ According to [11], Proposition 3.8, ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ is commutative.
+Let $X_{0},X_{1},\cdots,X_{g}$ be the $g+1$ variables. We define the automorphisms
+$w_{j}:\mathbb{C}\big[X_{0}^{\pm 1},X_{1}^{\pm 1},\cdots,X_{g}^{\pm 1}\big]\longrightarrow\mathbb{C}\big[X_{0}^{\pm 1},X_{1}^{\pm 1},\cdots,X_{g}^{\pm 1}\big],\quad 1\leq j\leq g$
+by
+$w_{j}(X_{0})=X_{0}X_{j}^{-1},\ \ \ w_{j}(X_{j})=X_{j}^{-1},\ \ \ w_{j}(X_{k})=X_{k}\ \ \textrm{for}\ k\neq 0,j.$
+Let $W_{g}$ be the finite group generated by $w_{1},\cdots,w_{g}$ and the permutations of variables $X_{1},\cdots,X_{g}$. Obviously $w_{j}^{2}$ is the identity map and $|W_{g}|=2^{g}g!$.
+Theorem 4.4.
+There exists an isomorphism
+$Q:{\mathscr{H}}(\Gamma_{g},\Delta_{g,p})\longrightarrow\mathbb{C}\big[X_{0}^{\pm 1},X_{1}^{\pm 1},\cdots,X_{g}^{\pm 1}\big]^{W_{g}}.$
+In fact, $Q$ is defined by
+$Q\big(\sum_{j=1}^{h}\Gamma_{g}M_{j}\big)=\sum_{j=1}^{h}Q(\Gamma_{g}M_{j})=\sum_{j=1}^{h}X_{0}^{-k_{0}(j)}\prod_{\nu=1}^{g}\big(p^{-\nu}X^{\nu}\big)^{k_{\nu}(j)}|\det A_{j}|^{g+1},$
+where we chose the representative $M_{j}$ of $\Gamma_{g}M_{j}$ of the form
+$M_{j}=\begin{pmatrix}A_{j}&B_{j}\\ 0&D_{j}\end{pmatrix},\quad A_{j}=\begin{pmatrix}p^{k_{1}(j)}&\ldots&*\\ 0&\ddots&\vdots\\ 0&0&p^{k_{g}(j)}\end{pmatrix}$.
+We note that the integers $k_{1}(j),\cdots,k_{g}(j)$ are uniquely determined.
+Proof. The proof can be found in [25, pp. 254–261]. $\square$
+For a prime $p$, we let
+${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})_{\mathbb{Q}}:=\left\{\,\sum c_{j}\,\Gamma_{g}M_{j}\Gamma_{g}\in{\mathscr{H}}(\Gamma_{g},\Delta_{g,p})\,|\ c_{j}\in\mathbb{Q}\ \right\}$
+be the $\mathbb{Q}$-algebra contained in ${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$. We put
+$G_{p}:=GSp(g,\mathbb{Q}_{p})\quad\textrm{and}\quad K_{p}=GSp(g,\mathbb{Z}_{p}).$
+We can identify ${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})_{\mathbb{Q}}$ with the $\mathbb{Q}$-algebra ${\mathscr{H}}_{g,p}^{\mathbb{Q}}$ of $\mathbb{Q}$-valued locally constant, $K_{p}$-biinvariant functions on $G_{p}$ with compact support.
+The multiplication on ${\mathscr{H}}_{g,p}^{\mathbb{Q}}$ is given by
+$(f_{1}*f_{2})(h)=\int_{G_{p}}f_{1}(g)\,f_{2}(g^{-1}h)dg,\quad f_{1},f_{2}\in{\mathscr{H}}_{g,p}^{\mathbb{Q}},$
+where $dg$ is the unique Haar measure on $G_{p}$ such that the volume of $K$ is $1$.
+The correspondence is obtained by sending the double coset $\Gamma_{g}M\Gamma_{g}$ to the characteristic function of $K_{p}MK_{p}$.
+In order to describe the structure of ${\mathscr{H}}_{g,p}^{\mathbb{Q}}$, we need to understand the $p$-adic Hecke algebras of the diagonal torus $\mathbb{T}$} and the Levi subgroup $\mathbb{M}$} of the standard parabolic group.
+Indeed, $\mathbb{T}$} is defined to be the subgroup consisting of diagonal matrices in $\Delta_{g}$ and
+${\mathbb{M}}=\left\{\,\begin{pmatrix}A&0\\ 0&D\end{pmatrix}\in\Delta_{g}\ \right\}$
+is the Levi subgroup parabolic subgroup
+$\left\{\,\begin{pmatrix}A&B\\ 0&D\end{pmatrix}\in\Delta_{g}\ \right\}.$
+Let $Y$ be the co-character group of $\mathbb{T}$}, i.e., $Y=\textrm{Hom}({\mathbb{G}}_{m},{\mathbb{T}}).$ We define the local Hecke algebra ${\mathscr{H}}_{p}({\mathbb{T}})$ for $\mathbb{T}$} to be the $\mathbb{Q}$-algebra of $\mathbb{Q}$-valued, ${\mathbb{T}}(\mathbb{Z}_{p})$-biinvariant functions on $\mathbb{T}(\mathbb{Q}_{p})$ with compact support.
+Then ${\mathscr{H}}_{p}(\mathbb{T})\cong\mathbb{Q}[Y],$ where $\mathbb{Q}[Y]$ is the group algebra over $\mathbb{Q}$ of $Y$.
+An element $\lambda\in Y$ corresponds the characteristic function of the double coset $D_{\lambda}=K_{p}\lambda(p)K_{p}$.
+It is known that ${\mathscr{H}}_{p}({\mathbb{T}})$ is isomorphic to the ring $\mathbb{Q}\big[(u_{1}/v_{1})^{\pm 1},\cdots,(u_{g}/v_{g})^{\pm 1},(v_{1}\cdots v_{g})^{\pm 1}\big]$ under the map
+$(a_{1},\cdots,a_{g},c)\mapsto(u_{1}/v_{1})^{a_{1}}\cdots(u_{g}/v_{g})^{a_{g}}(v_{1}\cdots v_{g})^{c}.$
+Similarly we have a $p$-adic Hecke algebra ${\mathscr{H}}_{p}({\mathbb{M}})$.
+Let $W_{\Delta_{g}}=N(\mathbb{T})/\mathbb{T}$ be the Weyl group with respect to $(\mathbb{T},\Delta_{g})$, where $N(\mathbb{T})$ is the normalizer of $\mathbb{T}$ in $\Delta_{g}$.
+Then $W_{\Delta_{g}}\cong S_{g}\ltimes(\mathbb{Z}/2\mathbb{Z})^{g},$ where the generator of the $i$-th factor $\mathbb{Z}/2\mathbb{Z}$ acts on a matrix of the form $\textrm{diag}(a_{1},\cdots,a_{g},d_{1},\cdots,d_{g})$ by interchanging $a_{i}$ and $d_{i}$, and the symmetry group $S_{g}$ acts by permuting the $a_{i}$$s$ and $d_{i}$$s$.
+We note that $W_{\Delta_{g}}$ is isomorphic to $W_{g}$.
+The Weyl group $W_{\mathbb{M}}$ with respect to $(\mathbb{T},\mathbb{M})$ is isomorphic to $S_{g}$.
+We can prove that the algebra ${\mathscr{H}}_{p}({\mathbb{T}})^{W_{\Delta_{g}}}$ of $W_{\Delta_{g}}$-invariants in ${\mathscr{H}}_{p}({\mathbb{T}})$ is isomorphic to $\mathbb{Q}\big[Y_{0}^{\pm 1},Y_{1},\cdots,Y_{g}\big]$ (cf.
+[25]).
+We let
+$B=\left\{\,\begin{pmatrix}A&B\\ 0&D\end{pmatrix}\in\Delta_{g}\,\Big|\ A\ \textrm{is upper triangular,}\ D\ \textrm{is lower trianular}\ \right\}$
+be the Borel subgroup of $\Delta_{g}$. A set $\Phi^{+}$ of positive roots in the root system $\Phi$ determined by $B$. We set $\rho={\frac{1}{2}}\sum_{\alpha\in\Phi^{+}}\alpha.$
+Now we have the map $\alpha_{\mathbb{M}}:\mathbb{M}\longrightarrow{\mathbb{G}}_{m}$ defined by
+$\alpha_{\mathbb{M}}(M):=l(M)^{-{{g(g+1)}\over 2}}\big(\det A\big)^{g+1},\quad M=\begin{pmatrix}A&0\\ 0&D\end{pmatrix}\in\mathbb{M}$
+and the map $\beta_{\mathbb{T}}:\mathbb{T}\longrightarrow{\mathbb{G}}_{m}$ defined by
+$\beta_{\mathbb{T}}(\textrm{diag}(a_{1},\cdots,a_{g},d_{1},\cdots,d_{g})):=\prod_{i=1}^{g}a_{1}^{g+1-2i},\quad\textrm{diag}(a_{1},\cdots,a_{g},d_{1},\cdots,d_{g})\in\mathbb{T}.$
+Let $\theta_{\mathbb{T}}:=\alpha_{\mathbb{M}}\,\beta_{\mathbb{T}}$ be the character of $\mathbb{T}.$ The $\textit{Satake's spherical map}\ S_{p,\mathbb{M}}:{\mathscr{H}}_{g,p}^{\mathbb{Q}}\longrightarrow{\mathscr{H}}_{p}(\mathbb{M})$ is defined by
+(4.4) $S_{p,\mathbb{M}}(\phi)(m):=|\alpha_{\mathbb{M}}(m)|_{p}\int_{U(\mathbb{Q}_{p})}\phi(mu)du,\quad\phi\in{\mathscr{H}}_{g,p}^{\mathbb{Q}},\ m\in\mathbb{M},$
+where $|\ \ |_{p}$ is the $p$-adic norm and $U(\mathbb{Q}_{p})$ denotes the unipotent radical of $\Delta_{g}$.
+Also another $\textit{Satake's spherical map}\ S_{\mathbb{M},\mathbb{T}}:{\mathscr{H}}_{p}({\mathbb{M}})\longrightarrow{\mathscr{H}}_{p}(\mathbb{T})$ is defined by
+(4.5) $S_{\mathbb{M},\mathbb{T}}(f)(t):=|\beta_{\mathbb{T}}(t)|_{p}\int_{\mathbb{M}\cap{\mathbb{N}}}f(tn)dn,\quad t\in{\mathscr{H}}_{p}(\mathbb{T}),\ t\in\mathbb{T},$
+where $\mathbb{N}$} is a nilpotent subgroup of $\Delta_{g}$.
+Theorem 4.5.
+The Satake’s spherical maps $S_{p,\mathbb{M}}$ and $S_{\mathbb{M},\mathbb{T}}$ define the isomorphisms of $\mathbb{Q}$-algebras
+(4.6) ${\mathscr{H}}_{g,p}^{\mathbb{Q}}\cong{\mathscr{H}}_{p}(\mathbb{T})^{W_{\Delta_{g}}}\quad\textrm{and}\quad{\mathscr{H}}_{p}(\mathbb{M})\cong{\mathscr{H}}_{p}(\mathbb{T})^{W_{\mathbb{M}}}.$
+We define the elements $\phi_{k}\,(0\leq k\leq g)$ in ${\mathscr{H}}_{p}(\mathbb{M})$ by
+$\phi_{k}:=p^{-{{k(k+1)}\over 2}}\,\mathbb{M}(\mathbb{Z}_{p})\begin{pmatrix}I_{g-k}&0&0\\ 0&pI_{g}&0\\ 0&0&I_{k}\end{pmatrix}\mathbb{M}(\mathbb{Z}_{p}),\quad i=0,1,\cdots,g.$
+Then we have the relation
+(4.7) $S_{p,\mathbb{M}}(T(p))=\sum_{k=0}^{g}\phi_{k}$
+and
+(4.8) $S_{p,\mathbb{M}}\big(T_{i}(p^{2})\big)=\sum_{j,k\geq 0,\,i+j\leq k}m_{k-j}(i)\,p^{-{k-j+1\choose 2}}\phi_{j}\phi_{k},$
+where
+$m_{s}(i):=\sharp\left\{\,A\in M(s,{\mathbb{F}}_{p})\,|\ {}^{t}A=A,\quad\textrm{corank}(A)=i\ \right\}.$
+Moreover, for $k=0,1,\cdots,g$, we have
+(4.9) $S_{\mathbb{M},\mathbb{T}}(\phi_{k})=(v_{1}\cdots v_{g})E_{k}(u_{1}/v_{1},\cdots,u_{g}/v_{g}),$
+where $E_{k}$ denotes the elementary symmetric function of degree $k$. The proof of (4.7)–(4.9) can be found in [28, pp. 142–145].
+4.2. Action of the Hecke Algebra on Siegel Modular Forms
+Let $(\rho,V_{\rho})$ be a finite dimensional irreducible representation of $GL(g,\mathbb{C})$ with highest weight $(k_{1},\cdots,k_{g})$.
+For a function $F:\mathbb{H}_{g}\longrightarrow V_{\rho}$ and $M\in\Delta_{g}^{+},$ we define
+$(f|_{\rho}M)(\Omega)=\rho(C\Omega+D)^{-1}f(M\cdot\Omega),\quad M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\Delta_{g}^{+}.$
+It is easily checked that $f|_{\rho}M_{1}M_{2}=\big(f|_{\rho}M_{1}\big)|_{\rho}M_{2}$ for $M_{1},M_{2}\in\Delta_{g}^{+}.$
+We now consider a subset $\mathscr{M}$} of $\Delta_{g}$ satisfying the following properties (M1) and (M2) :
+(M1) ${\mathscr{M}}=\bigcup_{j=1}^{h}\Gamma_{g}M_{j}\quad$ (disjoint union);
+(M2) ${\mathscr{M}}\,\Gamma_{g}\subset{\mathscr{M}}.$
+For a Siegel modular form $f\in M_{\rho}(\Gamma_{g})$, we define
+(4.10) $T({\mathscr{M}})f:=\sum_{j=1}^{h}f|_{\rho}M_{j}.$
+This is well defined, i.e., is independent of the choice of representatives $M_{j}$ because of the condition (M1).
+On the other hand, it follows from the condition (M2) that $T({\mathscr{M}})f|_{\rho}\gamma=T({\mathscr{M}})f$ for all $\gamma\in\Gamma_{g}.$ Thus we get a linear operator
+(4.11) $T({\mathscr{M}}):M_{\rho}(\Gamma_{g})\longrightarrow M_{\rho}(\Gamma_{g}).$
+We know that each double coset $\Gamma_{g}M\Gamma_{g}$ with $M\in\Delta_{g}$ satisfies the condition $(M1)$ and $(M2)$.
+Thus a linear operator $T({\mathscr{M}})$ defined in Formula (4.10) induces naturally the action of the Hecke algebra ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$ on $M_{\rho}(\Gamma_{g})$.
+More precisely, if ${\mathscr{N}}=\sum_{j=1}^{h}c_{j}\Gamma_{g}M_{j}\Gamma_{g}\in{\mathscr{H}}(\Gamma_{g},\Delta_{g})$, we define
+$T({\mathscr{N}})=\sum_{j=1}^{h}c_{j}T(\Gamma_{g}M_{j}\Gamma_{g}).$
+Then $T({\mathscr{N}})$ is an endomorphism of $M_{\rho}(\Gamma_{g})$.
+Now we fix a Siegel modular form $F$ in $M_{\rho}(\Gamma_{g})$ which is an eigenform of the Hecke algebra ${\mathscr{H}}(\Gamma_{g},\Delta_{g})$.
+Then we obtain an algebra homomorphism $\lambda_{F}:{\mathscr{H}}(\Gamma_{g},\Delta_{g})\longrightarrow\mathbb{C}$ determined by
+$T(F)=\lambda_{F}(T)F,\quad T\in{\mathscr{H}}(\Gamma_{g},\Delta_{g}).$
+By Theorem 4.4 or Theorem 4.5, one has
+${\mathscr{H}}(\Gamma_{g},\Delta_{g,p})$ $\cong$ ${\mathscr{H}}_{g,p}^{\mathbb{Q}}\otimes\mathbb{C}\cong\mathbb{C}[Y]^{W_{g}}$
+$\cong$ ${\mathscr{H}}_{p}({\mathbb{T}})^{W_{g}}\otimes\mathbb{C}$
+$\cong$ $\mathbb{C}\big[(u_{1}/v_{1})^{\pm 1},\cdots,(u_{g}/v_{g})^{\pm 1},(v_{1}\cdots v_{g})^{\pm 1}\big]^{W_{g}}$
+$\cong$ $\mathbb{C}[Y_{0},Y_{0}^{-1},Y_{1},\cdots,Y_{g}],$
+where $Y_{0},Y_{1},\cdots,Y_{g}$ are algebraically independent. Therefore one obtains an isomorphism
+$\textrm{Hom}_{\mathbb{C}}\big({\mathscr{H}}(\Gamma_{g},\Delta_{g,p}),\mathbb{C}\big)\cong\textrm{Hom}_{\mathbb{C}}\big({\mathscr{H}}_{g,p}^{\mathbb{Q}}\otimes\mathbb{C},\mathbb{C}\big)\cong(\mathbb{C}^{\times})^{(g+1)}/W_{g}.$
+The algebra homomorphism $\lambda_{F}\in\textrm{Hom}_{\mathbb{C}}\big({\mathscr{H}}(\Gamma_{g},\Delta_{g,p}),\mathbb{C}\big)$ is determined by the $W_{g}$-orbit of a certain $(g+1)$-tuple $\big(\alpha_{F,0},\alpha_{F,1},\cdots,\alpha_{F,g}\big)$ of nonzero complex numbers, called the $p$-Satake parameters of $F$.
+For brevity, we put $\alpha_{i}=\alpha_{F,i},\ i=0,1,\cdots,g$.
+Therefore $\alpha_{i}$ is the image of $u_{i}/v_{i}$ and $\alpha_{0}$ is the image of $v_{1}\cdots v_{g}$ under the map $\Theta$.
+Each generator $w_{i}\in W_{\Delta_{g}}\cong W_{g}$ acts by
+$w_{j}(\alpha_{0})=\alpha_{0}\alpha_{j}^{-1}\quad w_{j}(\alpha_{j})=\alpha_{j}^{-1},\quad w_{j}(\alpha_{k})=0\ \textrm{if}\ k\neq 0,j.$
+These $p$-Satake parameters $\alpha_{0},\alpha_{1}\cdots,\alpha_{g}$ satisfy the relation
+$\alpha_{0}^{2}\alpha_{1}\cdots\alpha_{g}=p^{\sum_{i=1}^{g}k_{i}-g(g+1)/2}.$
+Formula (4.12) follows from the fact that $T_{g}(p^{2})=\Gamma_{g}(pI_{2g})\Gamma_{g}$ is mapped to
+$p^{-g(g+1)/2}\,(v_{1}\cdots v_{g})^{2}\prod_{i=1}^{g}(u_{i}/v_{i}).$
+We refer to [25, p.
+258] for more detail.
+According to Formulas (4.7)–(4.9), the eigenvalues $\lambda_{F}\big(T(p)\big)$ and $\lambda_{F}\big(T_{i}(p^{2})\big)$ with $1\leq i\leq g$ are given respectively by
+(4.12) $\lambda_{F}\big(T(p)\big)=\alpha_{0}(1+E_{1}+E_{2}+\cdots+E_{g})$
+and
+(4.13) $\lambda_{F}\big(T_{i}(p^{2})\big)=\sum_{j,k\geq 0,\,j+i\leq k}^{g}m_{k-j}(i)\,p^{-{{k-j+1}\choose 2}}\,\alpha_{0}^{2}E_{j}E_{k},\quad i=1,\cdots,g,$
+where $E_{j}$ denotes the elementary symmetric function of degree $j$ in the variables $\alpha_{1},\cdots,\alpha_{g}$.
+The point is that the above eigenvalues $\lambda_{F}\big(T(p)\big)$ and $\lambda_{F}\big(T_{i}(p^{2})\big)\ (1\leq i\leq g)$ are described in terms of the $p$-Satake parameters $\alpha_{0},\alpha_{1}\cdots,\alpha_{g}$.
+Example 4.6.
+Suppose $g(\tau)=\sum_{n\geq 1}a(n)\,e^{2\pi in\tau}$ is a normalized eigenform in $S_{k}(\Gamma_{1})$. Let $p$ be a prime. Let $\beta$ be a complex number determined by the relation
+$(1-\beta X)(1-{\bar{\beta}}X)=1-a(p)X+p^{k-1}X^{2}.$
+Then
+$\beta+{\bar{\beta}}=a(p)\quad\textrm{and}\quad\beta{\bar{\beta}}=p^{k-1}.$
+The $p$-Satake parameters $\alpha_{0}$ and $\alpha_{1}$ are given by
+$(\alpha_{0},\alpha_{1})=\left(\beta,{{\bar{\beta}}\over\beta}\right)\quad or\quad\left({\bar{\beta}},{{\beta}\over{\bar{\beta}}}\right).$
+It is easily checked that $\alpha_{0}^{2}\alpha_{1}=\beta{\bar{\beta}}=p^{k-1}$ (cf. Formula (4.12)).
+Example 4.7.
+For a positive integer $k$ with $k>g+1$, we let
+$G_{k}(\Omega):=\sum_{M\in\Gamma_{g,0}\backslash\Gamma_{g}}\det(C\Omega+D)^{k},\quad M=\begin{pmatrix}A&B\\ C&D\end{pmatrix}$
+be the Siegel Eisenstein series of weight $k$ in $M_{k}(\Gamma_{g})$, where
+$\Gamma_{g,0}:=\left\{\begin{pmatrix}A&B\\ 0&D\end{pmatrix}\in\Gamma_{g}\right\}$
+is a parabolic subgroup of $\Gamma_{g}$.
+It is known that $G_{k}$ is an eigenform of all the Hecke operators (cf.
+[25, p.
+268]).
+Let $S_{1},\cdots,S_{h}$ be a complete system of representatives of positive definite even unimodular integral matrices of degree $2k$.
+We define the theta series $\theta_{S_{\nu}}$ by
+$\theta_{S_{\nu}}(\Omega):=\sum_{A\in\mathbb{Z}^{(2k,g)}}e^{\pi\,i\,\rm{Tr}(S_{\nu}[A]\Omega)},\quad 1\leq\nu\leq h.$
+If $k>g+1$, the Eisenstein series $G_{k}$ can be expressed as the weighted mean of theta series $\theta_{S_{1}},\cdots,\theta_{S_{h}}\,$:
+(4.14) $G_{k}(\Omega)=\sum_{\nu=1}^{h}m_{\nu}\,\theta_{S_{\nu}}(\Omega),\quad\Omega\in\mathbb{H}_{g},$
+where
+$m_{\nu}={{A(S_{\nu},S_{\nu})^{-1}}\over{A(S_{1},S_{1})^{-1}+\cdots+A(S_{h},S_{h})^{-1}}},\quad 1\leq\nu\leq h.$
+We recall that for two symmetric integral matrices $S$ of degree $m$ and $T$ of degree $n$, $A(S,T)$ is defined by
+$A(S,T):=\sharp\big\{\,G\in\mathbb{Z}^{(m,n)}\,|\ S[G]=\,{}^{t}GSG=T\,\big\}.$
+Formula (4.14) was obtained by Witt [29] as a special case of the analytic version of Siegel’s Hauptsatz.